Constructing and Assessing Income Statements Using Cost-to-Cost Method
Assume General Electric Company agreed in May 2016 to construct a nuclear generator for NSTAR, a utility company serving the Boston area. General Electric Company estimated that its construction costs would be $840 million. The contract price of $1,050 million is to be paid as follows: $350 million at the time of signing; $350 million on December 31, 2016; and $350 million at completion in May 2017. General Electric incurred the following costs in constructing the generator: $336 million in 2016 and $504 million in 2017.


a. Compute the amount of General Electric's revenue, expense, and income for both 2016 and 2017, and for both years combined, under the cost-to-cost revenue recognition method.
Enter dollar amounts in millions.

Cost-to-Cost Method

Year

Costs

incurred

% of total

excepted

costs

Revenue

recognized

Income

2016
2017
Total
Answer #1
Answer
Explanation :
Cost-to-Cost Method
Year Costs incurred % Total cost excepted costs Revenue recognized Income
2016 $              336 40% (336/840) $420 ($1050 x 40%) $                  84
2017 $              504 60% (504/840) $630 ($1050x60%) $                126
Total $              840 100% $                               1,050 $                210
Income in 2016 = 420-336 = 84
Income in 2017 = 630-504 = 126